point) A cashier at the Hungry Bear Restaurant in Disneyland's Critter Country serves an average of 25.5 customers per hour (believed to follow a Poisson distribution). Always conscious of managing park guests' waiting line experience, the disney operations research team wants to post a sign telling customers about the restaurant's efficiency. If the manager wants to be 85 % sure she doesn't overestimate, what time should she list? (enter answer to two decimal places)

Time per order = ......
Answer #1
here for exponential distribution parameter β =60/2.35 = 2.353 minute/order
pth percentile =-β*ln(1-p)
therefore 85th percentile =-2.353*ln(1-85/100)= 4.46 minutes per order